# Analyzing Character Representation in Media Content using Multimodal Foundation Model: Effectiveness and Trust

### v1

@@ Abstract.
Recent advances in AI has enabled automated analysis of complex media content at scale and generate actionable insights regarding character representation along such dimensions as gender and age. Past work focused on quantifying representation from audio/video/text using various ML models, but without having the audience in the loop. We ask, even if character distribution along demographic dimensions are available, how useful are they to the general public? Do they actually trust the numbers generated by AI models? Our work addresses these questions through a user study, while proposing a new AI-based character representation and visualization tool. Our tool based on the Contrastive Language Image Pretraining (CLIP) foundation model to analyze visual screen data to quantify character representation across dimensions of age and gender. We also designed effective visualizations suitable for presenting such analytics to lay audience. Next, we conducted a user study to seek empirical evidence on the usefulness and trustworthiness of the AI-generated results for carefully chosen movies presented in the form of our visualizations. We note that participants were able to understand the analytics from our visualization, and deemed the tool ‘overall useful’. Participants also indicated a need for more detailed visualizations to include more demographic categories and contextual information of the characters. Participants’ trust in AI-based gender and age models is seen to be moderate to low, although they were not against the use of AI in this context. Our tool including code, benchmarking, and data from the user study can be found here: https://anonymous.4open.science/r/Character-Representation-Media-FF7B
Multimodal foundation model, content analysis, gender and age representation, visualization, trust, empirical research
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Visualization design and evaluation methodsccs: Human-centered computing Visual analyticsccs: Computing methodologies Computer visionccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligenceccs: Applied computing Sociologyccs: Applied computing Media arts
Figure 1. Our tool to analyze the demographic character representation in videos. It detects the faces in the video, generates the CLIP image embedding for each face, and gives this as an input to a logistic regression gender and age classifiers. The analytics are visualized at the intersection of gender and age including information about the confidence and bias of the models. We conducted a user study to evaluate our tool with human participants.
@@ 1. Introduction
The problem of quantifying representation in media has been addressed by media researchers using traditional methods from screen research that involves hand counting (Smith et al., 2019; Communication and Marketing Staff, 2019). On the computational front, this is achieved using video/audio/language processing combined with machine learning (ML) models (Guha et al., 2015; Mazières et al., 2021; Somandepalli et al., 2021; Bamman et al., 2024). Such AI-based content analyses have been shown to uncover hidden biases in character portrayal in Hollywood movies, such as gendered gaps in screen and speaking time (Guha et al., 2015; Ramakrishna et al., 2017). The methodologies used and developed in the context of character representation have primarily focused on analysing the visual streams and faces (Bamman et al., 2024; Mazières et al., 2021). The general pipeline includes face detection followed by unsupervised or semi-supervised character discovery and classification of the extracted faces along various dimensions (Guha et al., 2015; Kulshreshtha and Guha, 2018; Bamman et al., 2024). A major challenge however is the variability and the long form of media. The characters’ appearances are diverse and may change widely in long form media (e.g., movies), which makes generalization difficult for ML models. This is where foundation models are expected to play an important role with their superior ability to generalize across diverse content. This current paper demonstrates the effective use of the multimodal foundation model, Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021), to address the task of quantifying character representation in media content.
@@
In this paper, we explore the effectiveness of the vision-language foundation model, CLIP (Radford et al., 2021), for analyzing the demographic dimensions (gender and age) of the characters’ faces that appear on screen. Following CLIP’s zero-shot paradigm, we train two classifiers to predict perceived gender and age; the models use CLIP embeddings of the face images detected in long videos. Our gender and age detectors are trained and benchmarked on a large publicly available dataset before they are used to extract analytics from the full-length movies. Benchmarking results show that our gender and age models perform close to the state-of-the-art. We extract analytics from three full length movies and create visualizations with the lay audience as target users. The design of the visualizations is informed by Our visualization is inspired by graphs commonly used to present demographic data to the public (Google, 2014; GDI et al., 2024; GDI, 2021a; Smith et al., 2019; GDI and Lyda Hill Foundation, 2018). Next, we provide empirical evidence on the usefulness and trustworthiness of the AI-generated results presented in form of our visualizations through a user study. The user study showed that the presented information was well understood, the tool was deemed overall useful, although participants would be interested in seeing more demographics, mainly ethnicity, and contextual characters’ analyses. Participants’ trust in our gender and age models was mediocre and low, respectively. To summarize, our main contributions are:
@@ 2. Related work
Traditionally, automated analysis of media content has been focused on addressing the needs of organizing, indexing and navigating through large media data corpora. More recently, computing effort is being driven towards generating insights and human-centred analytics from large volumes of media content across audio, video and text (Bamman et al., 2024; Mazières et al., 2021; Somandepalli et al., 2021). Unimodal approaches such as those of Bamman et al. (2024) and Mazières et al. (2021) automatically quantify character representation in media through analysing characters faces. These works used the visual data stream with classical ML models to analyse characters faces. Bamman et al. (2024) also studies on ethnic diversity, but this was done through human annotations instead of ML. Ramakrishna et al. (2017) combined linguistic analysis with ML to reveal systematic patterns in character portrayals, such as ethnic minorities not interacting with other characters.
Existing work on multimodal approaches leverage the inherent multimodal nature of media data to analyse content for understanding character representation. The earliest work include that of Guha et al. (2015) which developed a multimodal system involving speaker diarization (i.e., who speaks when), face detection and gender identification to measure speaking time and screen time of male and female actors. Researchers have also used cross-domain information to enhance basic tasks such as gender detection or active speaker detection in movies (Hebbar et al., 2018; Sharma et al., 2019). Recent advancements in multimodal foundation models and large language models (LLMs) have led to a shift of the content analysis towards this direction. Gan et al. (2023) develop a CLIP-based method leveraging text and image analysis for apparent personality perception. Cerit et al. (2025) developed the Media Content Atlas, which is a pipeline to explore media content in the multidimensional media space using multimodal LLMs.
@@ 3. Methodology
Fig. 1 presents an overview of the video analysis pipeline along with the visualization tool that we developed. Below we present the methodology used to develop this tool.
@@ 3.1. CLIP-based video analysis
Face detection and CLIP embeddings. Our video analysis pipeline, like several past work, relies on analyzing only the characters’ faces that are seen in the videos. We use the DeepFace’s (Serengil and Ozpinar, 2020) Single Shot Face Detector (Zhang et al., 2017) as our face detector. For each detected face we use the Huggingface transformers package (Wolf et al., 2019) to generate their CLIP embeddings. The CLIP model itself is trained on the WebImageText dataset (Radford et al., 2021).
@@
Table 1. Gender and age prediction performance of CLIP-based LR models on the FairFace dataset.
@@
Results on full-length films. We chose three full-length movies to generate the gender and age analytics to be used for visualization and subsequent user study. This include Mamma Mia! The Movie (2008), The Best Exotic Marigold Hotel (2011), and Black Panther (2018). The films were selected to include diverse distributions of on-screen appearances for gender and age. The results are shown in Table 2.
@@ 3.2. The visualization tool
Fig. 1 presents our visualization created to show the gender and age representation at their intersection i.e., for women characters who are over 50. Our visualization is inspired by the doughnut charts commonly used to present demographic data to the public (Google, 2014; GDI et al., 2024; GDI, 2021a; Smith et al., 2019; GDI and Lyda Hill Foundation, 2018). Such graphs are widely used in the news, reports, etc., and the general audience is expected to be familiar with them. We alter the standard design of the doughnut chart to embed a second ring within it with the purpose to present the distribution of appearances at the intersection of gender and age. The visualization is enhanced with popups on hover to show the exact percentages corresponding to each demographic group and explanations for what the AI prediction confidence and bias means.
In the case of age, we map the age predictions to the set {’Up to 50’,’Over 50’}. We assign each detected face to the appropriate group in this set depending on which of the following sums was greater: the sum of confidences for all age groups with ages¿50 and the corresponding sum for ages¡50.
@@ 4.1. Procedure
We generate the visualizations (see Fig. 2) for the 3 films named in Table 2 and use them to evaluate our tool with human users. We do not disclose the names of the films to the participants to avoid having any bias arising due to prior knowledge or perception. We recruited 30 participants through mailing lists and social media channels within our institution. The participants spent 30 minutes at max and they were compensated for their time. The user study was approved by the relevant institutional Ethics Review Committee. Only after the participants provided consent, they received a link to a questionnaire form and were able to proceed to the study.
@@
The user study comprised three parts: Part 1 is a pre-questionnaire capturing some participant demographics (questions noted as Q1.x). Fig. 3 presents the distribution of participants’ responses. Part 2 provides participants with a link to view the visualizations for the three films (Fig. 2). The questions in this part test their understanding of the presented information and trust in the ML models (questions noted as Q2.x). Part 3 consists of a post-questionnaire investigating the usefulness of our tool and the user experience (questions noted as Q3.x). Table 3 presents an overview of the question in Parts 2 and 3 mapped to the RQs.
@@ 4.2. Analysis
We conduct statistical Bayesian inference for the closed-ended questions similarly to Taka et al. (2024). These questions can be categorized in two groups for the purpose of analysis.
First, questions providing a list of options to select the correct response(s) (Q2.1-8 - understandability). We measure the accuracy of each participant in every question. Participants’ responses are transformed into a binary representation with 0 indicating a wrong and 1 a correct selection. Participants that selected the “I don’t know” (idk) option were excluded. Participants’ transformed responses in Q2.1-7 (single selection was allowed) consist of a single digit. In Q2.8 (multiple selections were allowed) this number was 4 (excluding the idk option). Participants’ performance in each question is calculated as the number of occurrences of digit 1 in their response. We assume a Bernoulli/binomial likelihood for the number of correct responses (depending on whether the single or multiple selections were allowed) with a Beta prior for the probability of a correct answer. Then, we estimate the posterior distribution of this probability as a measure of the users’ understanding of the presented information.
Second, questions asking for a score in the range of a Likert scale (Q2.9-10 - trust, Q3.1-5 - usefulness, Q3.6-9 - user experience), we assume a normal likelihood for the users’ scores with a normal prior for the mean and a half-normal prior for the standard deviation of the scores. We estimate the posterior distribution of the mean score as a measure of the user reported trust, usefulness, and user experience.
@@ 4.3. Results
Fig. 4a presents the Bayesian confidence intervals (CIs) for the probability of participants giving correct responses to questions Q2.1-8 based on the posterior distribution. A probability of $0.5$ indicates random responses. CIs that are on the right of the reference line of $0.5$ and do not intersect with it indicate the probability of providing accurate responses. The probability of correct response increases with the distance of the CI from the reference line on the right side of it. Fig. 4b presents the (Bayesian) CIs for the mean score participants provided in questions Q2.9-10 and Q3.1-9.
@@
Information understandability (RQ1). Based on our results, the evidence on the ability of the participants to understand the character distribution the doughnut charts’ present (cast, characters, on-screen appearances etc.) is not strong (Q2.1).The CI intersects with the reference line and is quite wide spanning between probability 0.47 and 0.78 (mean=0.62). Although the title of the visualizations clearly states that the distribution of on-screen appearances is presented (Fig. 2), only 19 out of 30 participants noted this correctly, while 7 participants assumed the distribution refers to the cast of the film, and 4 to the number of characters appearing in the films. Our results suggest that participants are able to distinguish gender and age information. The participants can identify which categories of appearances between Female/Male or Over/Up to 50 prevail in the films. The CIs for the corresponding questions (Q2.2-3) indicate a high probability of correct responses, above $0.75$ with mean probability $0.87$ and $0.86$, respectively. The design of the visualization has been crucial here: the required information to respond to these questions is clearly stated in the center of the doughnut charts, and is indicated by the colored wedges. This argument is further supported by the results in Q2.5. Participants are able to correctly identify the film with the fair representation of the category “Over 50”, albeit the wider confidence interval for the probability of correct response (it spans between 0.6 and 0.89, mean=0.75).
@@
We also investigated participants’ ability to interpret the information presented at the intersection of gender and age. Responses to Q2.4 were ambiguous, and did not provide us with reliable evidence. This question asked which category at the intersection of gender and age appears the least in Film 3, while “Female Over 50” does not appear at all and “Male Over 50” appears the least. We considered the first as a correct response, but the second could be deemed correct, as well. In practice, 18 participants selected the first, while 7 selected the second and another 3 selected the “Male Up to 50” and 2 the “Female Up to 50”. The CI for Q2.4 intersects with the reference line and is quite wide (it spans between 0.43 and 0.75, mean=0.59).
@@
Nevertheless, Q2.6 provides a good evidence for this. Participants correctly identify the film that proportionately favors the representation of the category “Female Over 50”, albeit the wider CI for the probability of correct response (it spans between 0.64 and 0.91, mean=0.78). The design of the visualization about this point is more complex: the inner circle of the doughnut charts presented the percentage of representation for the groups at the intersection of gender and age with appropriate splits in wedges. The grayed wedges had to be considered here. Interactive pop-ups showed the label of each wedge (colored or not) and the associated percentage.
@@
Finally, our results also suggest that participants are able to identify the film with the greatest AI confidence for age prediction (Q2.7) and the category that the AI could likely under-detect (Q2.8). We found a quite wide CI for Q2.7 with probability between 0.61 and 0.89 (mean=0.75), but a much shorter CI for Q2.8 between 0.72 and 0.86 (mean=0.79). The design to present this information was quite straightforward. The AI prediction confidence was provided as text below the doughnut charts. The AI bias was provided in the form of bar graphs.
@@
The CI for the usefulness of the AI use in this context spans between 2.98 and 3.77 (mean=3.39). This implies that participants admit some usefulness of the AI in this case and do not reject it.
@@
Three participants mentioned that they would be interested to see more demographics like ethnicity/race in this tool (“Develop it further to recognise race, language etc.”,“I would be interested to see how accurate AI can analyse the gender and age of people from different ethnic background.”,“To check for cultural bias”).
Eight participants mentioned that they would (probably) not use this tool, and another 2 were not sure about its applicability. Two of them implied that the counts of appearances is not enough to account for how characters are represented (“Representation is very important, but I’m not convinced that the best way to get there is through numerical comparisons.”, “There are so many ways this data could be skewed and just having data on gender and age doesn’t mean a film reflects positively on the demographics represented. For example, movies like Seven Brides for Seven Brothers would have high instances of women on screen, but that doesn’t mean the depicitons of women are positive, so to me, a tool like this has absolutely no practical use without further detail provided. Or like in horror movies–yes, there may be many women on screen, but are they all there because they’re being tortured and murdered? A tool like this ignores the context in which people appear on screen and would end up generating a lot of false positives in terms of contemporary metrics of representation–it’s not enough to just count how many women appear on screen anymore. We need to consider *how* they’re depicted.”). Another participant also mentioned that they would use such a tool to study “shifts in representation over time. Expand it to identify shades of characters (e.g. hero/villain, positive/negative roles, victim/perpetrator etc.)”. One of these participants also expressed their skepticism towards the use of AI. Four people stated that they are not interested in the presented information with 2 of them mentioning that they are primarily interested in the cast of the films.
@@
$\bullet$“The tool seems useful, but only if I want to filter movies based on gender” $\bullet$ “I like that the parts that are not relevant to subject […] are greyed out.” $\bullet$“I like the visualisation idea a lot” $\bullet$“Information was presented in a very neat and clean manner” $\bullet$ “I think it is a cool tool, and it’s impressive how accurate it seems to be” $\bullet$ “The graphs were easy to understand” $\bullet$ “I like the te[c]h data is presented with pie charts” $\bullet$ “This tool was ok-ish I guess. Good observations to be honest.” $\bullet$ “I liked the pop ups next to the confidence and bias headings, that was needed.” $\bullet$ “I also really liked the info icon for understanding some of the terms” $\bullet$ “It was extremely useful to see the comparative values for actual and predicted.” $\bullet$ “I liked the figures below the pie chart” $\bullet$ “the hover features definitely made it a lot easier to interpret so that was helpful”.
Ten participants found that the presentation was confusing or hard-to-read or understand. Specifically, regarding the amount of presented information, 1 participant mentioned that too much information was presented, 2 participants mentioned a cognitive overload by the number of films/charts presented, and 2 participants also mentioned that they needed some time to understand the graphs. Regarding the design of the visualization, 2 participants stated that the doughnut/pie charts were not appropriate in this context, 3 stated that the encasing of the rings for the gender and age was confusing for them, 3 mentioned that the color coding was confusing (e.g., “it would be easier to understand if there were different colours for different categories, like Male over 50s is a different colour than Female over 50s”, “green represents male and female aged over 50 at the same time”), 2 did not like hovering to see the labels (another participant found this feature useful).
Another source of confusion mentioned by 3 participants was what the percentages represented/how they were calculated (e.g., “I am confused with how the percentages are calculated, whether it is counting the cast or measuring screentime.”. “[…] like if a character leaves the scene and then comes back does that count as two?”). Finally, 2 participants misunderstood (elements of) the design (e.g., “There was no information on male appearances”,“the description of male or females over 50 should’ve been specified”).

### v2

@@ Abstract.
Recent advances in AI has made automated analysis of complex media content at scale possible while generating actionable insights regarding character representation along such dimensions as gender and age. Past works focused on quantifying representation from audio/video/text using AI models, but without having the audience in the loop. We ask, even if character distribution along demographic dimensions are available, how useful are those to the general public? Do they actually trust the numbers generated by AI models? Our work addresses these open questions by proposing a new AI-based character representation tool and performing a thorough user study. Our tool has two components: (i) An analytics extraction model based on the Contrastive Language Image Pretraining (CLIP) foundation model that analyzes visual screen data to quantify character representation across age and gender; (ii) A visualization component effectively designed for presenting the analytics to lay audience. The user study seeks empirical evidence on the usefulness and trustworthiness of the AI-generated results for carefully chosen movies presented in the form of our visualizations. We found that participants were able to understand the analytics in our visualizations, and deemed the tool ‘overall useful’. Participants also indicated a need for more detailed visualizations to include more demographic categories and contextual information of the characters. Participants’ trust in AI-based gender and age models is seen to be moderate to low, although they were not against the use of AI in this context. Our tool including code, benchmarking, and the user study data can be found at https://github.com/debadyuti0510/Character-Representation-Media.
Multimodal foundation model, media content analysis, gender and age representation, visualization, AI trust
journalyear: 2025copyright: ccconference: Proceedings of the 27th International Conference on Multimodal Interaction; October 13–17, 2025; Canberra, ACT, Australiabooktitle: Proceedings of the 27th International Conference on Multimodal Interaction (ICMI ’25), October 13–17, 2025, Canberra, ACT, Australiadoi: 10.1145/3716553.3750785isbn: 979-8-4007-1499-3/2025/10ccs: Human-centered computing Visualization design and evaluation methodsccs: Human-centered computing Visual analyticsccs: Computing methodologies Computer visionccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligenceccs: Applied computing Sociologyccs: Applied computing Media arts
Figure 1. Our tool to analyze the demographic character representation in videos. It detects the faces in the video, generates the face CLIP embeddings and uses them as input to gender and age classifiers. The analytics are visualized at the intersection of gender and age including information about the confidence and bias of the models. We evaluated our tool with human users.
@@ 1. Introduction
The problem of quantifying representation in media has been addressed by media researchers using traditional methods from screen research that involves hand counting (Smith et al., 2019; Communication and Marketing Staff, 2019). On the computational front, this is achieved using video/audio/language processing combined with machine learning (ML) models (Guha et al., 2015; Mazières et al., 2021; Somandepalli et al., 2021; Bamman et al., 2024). Such AI-based content analyzes have been shown to uncover hidden biases in character portrayal in Hollywood movies, such as gendered gaps in screen and speaking time (Guha et al., 2015; Ramakrishna et al., 2017). The methodologies used and developed in the context of character representation have primarily focused on analysing the visual streams and faces (Bamman et al., 2024; Mazières et al., 2021). The general pipeline includes face detection followed by unsupervised or semi-supervised character discovery and classification of the extracted faces along various dimensions (Guha et al., 2015; Kulshreshtha and Guha, 2018; Bamman et al., 2024). A major challenge however is the variability and the long form of media. The characters’ appearances are diverse and may change widely in long form media (e.g., movies), which makes generalization difficult for ML models. This is where foundation models are expected to play an important role with their superior ability to generalize across diverse content. This current paper demonstrates the effective use of the multimodal foundation model, Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021), to address the task of quantifying character representation in media content.
@@
In this paper, we explore the effectiveness of the vision-language foundation model, CLIP (Radford et al., 2021), for analyzing the demographic dimensions (gender and age) of the characters’ faces that appear on screen. Following CLIP’s zero-shot paradigm, we train two classifiers to predict perceived gender and age; the models use CLIP embeddings of the face images detected in long videos. Our gender and age detectors are trained and benchmarked on a large publicly available dataset before they are used to extract analytics from the full-length movies. Benchmarking results show that our gender and age models perform close to the state-of-the-art. We extract analytics from three full length movies and create visualizations with the lay audience as target users. The design of the visualizations is informed by graphs commonly used to present demographic data to the public (Google, 2014; GDI et al., 2024; GDI, 2021a; Smith et al., 2019; GDI and Lyda Hill Foundation, 2018). Next, we provide empirical evidence on the usefulness and trustworthiness of the AI-generated results presented in form of our visualizations through a user study. To summarize, our main contributions are:
@@ 2. Related work
Traditionally, automated analysis of media content has been focused on addressing the needs of organizing, indexing and navigating through large media data corpora. More recently, computing effort is being driven towards generating insights and human-centred analytics from large volumes of media content across audio, video and text (Bamman et al., 2024; Mazières et al., 2021; Somandepalli et al., 2021). Unimodal approaches such as those of Bamman et al. (2024) and Mazières et al. (2021) automatically quantify character representation in media through analysing characters’ faces. These works used the visual data stream with classical ML models to analyze characters’ faces. Bamman et al. (2024) also studies on ethnic diversity, but this was done through human annotations instead of ML. Ramakrishna et al. (2017) combined linguistic analysis with ML to reveal systematic patterns in character portrayals, such as ethnic minorities not interacting with other characters.
Existing work on multimodal approaches leverage the inherent multimodal nature of media data to analyze content for understanding character representation. The earliest work include that of Guha et al. (2015), which developed a multimodal system involving speaker diarization (i.e., who speaks when), face detection and gender identification to measure speaking time and screen time of male and female actors. Researchers have also used cross-domain information to enhance basic tasks such as gender detection or active speaker detection in movies (Hebbar et al., 2018; Sharma et al., 2019). Recent advancements in multimodal foundation models and large language models (LLMs) have led to a shift of the content analysis towards this direction. Gan et al. (2023) developed a CLIP-based method leveraging text and image analysis for apparent personality perception. Cerit et al. (2025) developed the Media Content Atlas, which is a pipeline to explore media content in the multidimensional media space using multimodal LLMs.
@@ 3. Methodology
Fig. 1 presents an overview of the video analysis pipeline along with the visualization tool that we developed. Below we explain the methodology used to develop this tool.
@@ 3.1. CLIP-based video analysis
Face detection and CLIP embeddings. Our video analysis pipeline, like several past works, relies on analyzing only the characters’ faces that are seen in the videos. We use the DeepFace’s (Serengil and Ozpinar, 2020) Single Shot Face Detector (Zhang et al., 2017) as our face detector. For each detected face we use the Huggingface transformers package (Wolf et al., 2019) to generate their CLIP embeddings. The CLIP model itself is trained on the WebImageText dataset (Radford et al., 2021).
@@
Table 1. Gender and age prediction performance of the CLIP-based models on the FairFace dataset.
@@
Results on full-length films. We chose three full-length films to generate the gender and age analytics to be used for visualization and subsequent user study. These include Mamma Mia! The Movie (2008), The Best Exotic Marigold Hotel (2011), and Black Panther (2018). The films were selected to include diverse distributions of on-screen appearances for gender and age. The films and analysis results are shown in Table 2.
@@ 3.2. The visualization tool
Fig. 1 presents our visualization created to show the gender and age representation at their intersection i.e., for women characters over 50. Our visualization is inspired by the doughnut charts commonly used to present demographic data to the public (Google, 2014; GDI et al., 2024; GDI, 2021a; Smith et al., 2019; GDI and Lyda Hill Foundation, 2018). Such graphs are widely used in the news, reports, etc., and the general audience is expected to be familiar with them. We alter the standard design of the doughnut chart to embed a second ring within it with the purpose to present the distribution of appearances at the intersection of gender and age. The visualization is enhanced with popups on hover to show the exact percentages corresponding to each demographic group and explanations for what the AI prediction confidence and bias means.
In the case of age, we map the age predictions to the set {‘Up to 50’,‘Over 50’}. We assign each detected face to the appropriate group in this set depending on which of the following sums is greater: the sum of confidences for all age groups with ages¿50 and the corresponding sum for ages¡50.
@@ 4.1. Procedure
We generate the visualizations (see Fig. 2) for the 3 films named in Table 2 and use them to evaluate our tool with human users. We do not disclose the names of the films to participants to avoid any bias due to prior knowledge or perception. We recruited 30 participants through University of Glasgow mailing lists and social media channels. The participants spent 30 minutes at max and they were compensated for their time. The user study was approved by the Ethics Committee of University of Glasgow. After the participants provide consent, they can proceed to the study questions.
@@
The user study comprise three parts: Part 1 captures participant demographics (questions noted as Q1.x). Fig. 3 presents the distribution of Part 1 responses. Part 2 provides a link to the film visualizations (Fig. 2) and tests participants’ understanding of the presented information and trust in the ML models (questions noted as Q2.x). Part 3 investigates the tool usefulness and the user experience (questions noted as Q3.x). Table 3 presents an overview of the question in Parts 2 and 3 mapped to the RQs.
@@ 4.2. Analysis
We conduct statistical Bayesian inference for the closed-ended questions similarly to Taka et al. (2024) (Bayesian analysis can provide useful inferences even for small sample sizes). These questions can be categorized in two groups for the purpose of analysis.
First, questions providing a list of options to select the correct response(s) (Q2.1-8 - understandability). We measure the accuracy of each participant in every question. Participants’ responses are transformed into a binary representation with 0 indicating a wrong and 1 a correct selection. Participants that selected the “I don’t know” (idk) option were excluded. This option was included to avoid noise by random responses. Very few people picked this option (and not in all questions); Q2.5/6: 1, Q2.2: 2, Q2.3: 3, Q2.8: 4. Participants’ transformed responses in Q2.1-7 (single selection was allowed) consist of a single digit. In Q2.8 (multiple selections were allowed) this number was 4 (excluding the idk option). Participants’ performance in each question is calculated as the number of digit 1 occurrences in their response. We assume a Bernoulli/binomial likelihood for the number of correct responses (depending on whether the single or multiple selections were allowed) with a Beta prior for the probability of a correct answer. Then, we estimate the posterior distribution of this probability as a measure of the users’ understanding of the presented information.
Second, questions asking for a rating in a Likert scale (Q2.9-10 - trust, Q3.1-5 - usefulness, Q3.6-9 - user experience). We assume a normal likelihood for the ratings with a normal prior for the mean and a half-normal prior for the standard deviation of them. We estimate the posterior distribution of the mean rating as a measure of the user reported trust, usefulness, and user experience.
@@ 4.3. Results
Fig. 4a presents the Bayesian confidence intervals (CIs) for the probability of participants giving correct responses to questions Q2.1-8 based on the posterior distribution. A probability of $0.5$ indicates random responses. CIs that are on the right of the reference line of $0.5$ and do not intersect with it indicate the probability of providing accurate responses. The probability of correct response increases with the distance of the CI from the reference line on the right side of it. Fig. 4b presents the (Bayesian) CIs for the mean participant ratings in questions Q2.9-10 and Q3.1-9.
@@
Information understandability (RQ1). Based on our results, the evidence on the ability of the participants to understand the character distribution the doughnut charts’ present (cast, characters, on-screen appearances etc.) is not strong (Q2.1).The CI intersects with the reference line and is quite wide spanning between probability 0.47 and 0.78 (mean=0.62). Although the title of the visualizations clearly states that it’s the distribution of on-screen appearances presented (Fig. 2), only 19 out of 30 participants noted this correctly, while 7 participants assumed the distribution refers to the cast of the film, and 4 to the number of characters appearing in the films. Our results suggest that participants are able to distinguish gender and age information. The participants can identify which categories of appearances between Female/Male or Over/Up to 50 prevail in the films. The CIs for the corresponding questions (Q2.2-3) indicate a high probability of correct responses, above $0.75$ with mean probability $0.87$ and $0.86$, respectively. The design of the visualization has been crucial here: the required information to respond to these questions is clearly stated in the center of the doughnut charts, and is indicated by the colored wedges. This argument is further supported by the results in Q2.5. Participants are able to correctly identify the film with the fair representation of the category ‘Over 50’, albeit the wider confidence interval for the probability of correct response (it spans between 0.6 and 0.89, mean=0.75).
@@
We also investigated participants’ ability to interpret the information presented at the intersection of gender and age. Responses to Q2.4 were ambiguous, and did not provide us with reliable evidence. This question asked which category at the intersection of gender and age appears the least in Film 3, while ‘Female Over 50’ does not appear at all and ‘Male Over 50’ appears the least. We considered the first as a correct response, but the second could be deemed correct, as well. In action, 18 participants selected the first, 7 selected the second, 3 selected the ‘Male Up to 50’ and 2 the ‘Female Up to 50’. The CI for Q2.4 intersects with the reference line and is quite wide (it spans between 0.43 and 0.75, mean=0.59).
@@
Nevertheless, Q2.6 provides a good evidence for this. Participants correctly identify the film that proportionately favors the representation of the category ‘Female Over 50’, albeit the wider CI for the probability of correct response (it spans between 0.64 and 0.91, mean=0.78). The design of the visualization about this point is more complex: the inner circle of the doughnut charts presented the percentage of representation for the groups at the intersection of gender and age with appropriate splits in wedges. The grayed wedges had to be considered here. Interactive pop-ups showed the label of each wedge (colored or not) and the associated percentage.
@@
Finally, our results suggest that participants are able to identify the film with the greatest AI confidence for age prediction (Q2.7) and the category that the AI could likely under-detect (Q2.8). We found a quite wide CI for Q2.7 with probability between 0.61 and 0.89 (mean=0.75), but a much shorter CI for Q2.8 between 0.72 and 0.86 (mean=0.79). The presentation of this information was quite straightforward. The AI prediction confidence was provided in text and the AI bias in bar graphs below the doughnut charts.
@@
The CI for the usefulness of AI in this context spans between 2.98 and 3.77 (mean=3.39). This implies that participants admit some usefulness of the AI in this case and do not reject it.
@@
3 participants mentioned that they would be interested in seeing more demographics like ethnicity/race in this tool: $\bullet$ “Develop it further to recognize race, language etc.” $\bullet$ “I would be interested to see how accurate AI can analyze the gender and age of people from different ethnic background.” $\bullet$ “To check for cultural bias”.
8 participants mentioned that they would (probably) not use this tool, and another 2 were not sure about its applicability. 2 of them implied that the counts of appearances is not enough to account for how characters are presented: $\bullet$ “Representation is very important, but I’m not convinced that the best way to get there is through numerical comparisons.” $\bullet$ “There are so many ways this data could be skewed and just having data on gender and age doesn’t mean a film reflects positively on the demographics represented. For example, movies like Seven Brides for Seven Brothers would have high instances of women on screen, but that doesn’t mean the depictions of women are positive, so to me, a tool like this has absolutely no practical use without further detail provided. Or like in horror movies–yes, there may be many women on screen, but are they all there because they’re being tortured and murdered? A tool like this ignores the context in which people appear on screen and would end up generating a lot of false positives in terms of contemporary metrics of representation–it’s not enough to just count how many women appear on screen anymore. We need to consider *how* they’re depicted.”.
A participant mentioned that would use such a tool to study “shifts in representation over time. Expand it to identify shades of characters (e.g. hero/villain, positive/negative roles, victim/perpetrator etc.)”. A participant expressed skepticism about the use of AI. 4 people stated they have no interest in the presented information with 2 of them mentioning they are primarily interested in the film cast.
@@
$\bullet$“The tool seems useful, but only if I want to filter movies based on gender” $\bullet$ “I like that the parts that are not relevant to subject […] are greyed out.” $\bullet$“I like the visualization idea a lot” $\bullet$“Information was presented in a very neat and clean manner” $\bullet$ “I think it is a cool tool, and it’s impressive how accurate it seems to be” $\bullet$ “The graphs were easy to understand” $\bullet$ “I like the te[c]h data is presented with pie charts” $\bullet$ “This tool was ok-ish I guess. Good observations to be honest.” $\bullet$ “I liked the pop ups next to the confidence and bias headings, that was needed.” $\bullet$ “I also really liked the info icon for understanding some of the terms” $\bullet$ “It was extremely useful to see the comparative values for actual and predicted.” $\bullet$ “I liked the figures below the pie chart” $\bullet$ “the hover features definitely made it a lot easier to interpret so that was helpful”.
Ten participants found that the presentation was confusing or hard-to-read or understand. Specifically, regarding the amount of presented information, 1 participant mentioned that too much information was presented, 2 participants mentioned a cognitive overload by the number of films/charts presented, and 2 participants also mentioned that they needed some time to understand the graphs. Regarding the design of the visualization, 2 participants stated that the doughnut/pie charts were not appropriate in this context, 3 stated that the encasing of the rings for the gender and age was confusing for them, 3 mentioned that the color coding was confusing (e.g., $\bullet$ “it would be easier to understand if there were different colors for different categories, like Male over 50s is a different color than Female over 50s” $\bullet$ “green represents male and female aged over 50 at the same time”), 2 did not like hovering to see the labels (another participant found this feature useful).
Another source of confusion mentioned by 3 participants was what the percentages represented/how they were calculated (e.g., $\bullet$ “I am confused with how the percentages are calculated, whether it is counting the cast or measuring screentime.” $\bullet$ “[…] like if a character leaves the scene and then comes back does that count as two?”). Finally, 2 participants misunderstood (elements of) the design (e.g., $\bullet$ “There was no information on male appearances” $\bullet$ “the description of male or females over 50 should’ve been specified”).